OpenStudy (secret-ninja):

Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = - 1 over 5(x - 2) y - 2 = - 1 over 5(x + 3)

OpenStudy (bibby):

point slope form: \(y-y_1=m(x-x_1)\) what is m?

OpenStudy (secret-ninja):

- 0.2 right?

OpenStudy (bibby):

slope = \(\large \frac{y_2-y_1}{x_2-x_1}=\frac{1-2}{2-(-3)}\)

OpenStudy (secret-ninja):

Oh ok

OpenStudy (secret-ninja):

So it is negative 1 over 5 right?

OpenStudy (bibby):

yeah. now we've narrowed our choices down to 2. plug in to the equation

OpenStudy (secret-ninja):

Okay wait.. I think i did my work right so it would be D right? Y – 2 = -1over5(x+3)

OpenStudy (bibby):

yus

OpenStudy (secret-ninja):

Thank you!!